The percentage of residents in a city in Georgia who recycle increased from 28.3% in 2000 to 31.1% in 2007. Predict the percentage of residents who will recycle in 2017.
Answer by stanbon(75887)

The percentage of residents in a city in Georgia who recycle increased from 28.3% in 2000 to 31.1% in 2007. Predict the percentage of residents who will recycle in 2017.
----
You have 2 points (year,percent): (0,28.3) and (7,31.1)
---
slope = (31.1-28.3)/(7-0) = 0.4
---
intercept is 28.3
-----
Equation:
P(t) = 0.4t + 28.3 where t is the # of years after 2000
---
P(17) = 0.4*17+28.3 = 35.1%
=================================
Cheers,
Stan H.
=========